OpenStudy (anonymous):

what number would complete the square? x^2-2x+6=0

OpenStudy (anonymous):

a)1 b)6 c)4 d)-1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

@03453660

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

want to know??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

@03453660

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your equation is x^2-2x+6=0 now take the constant to RHS x^2-2x = -6 look to LHS of this equation what should i add to make it perfect square like it become (a+b)^2 or (a-b)^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so if i add 1 to LHS and RHS then LHS become a perfect square x^2-2x+1 = -6+1 (x-1)^2 = -5 this LHS has became a perfect square by adding 1 to it

OpenStudy (anonymous):

got it!! thanks!!

OpenStudy (anonymous):

welcome! and sorry for the late reply i have some problem in internet connection

OpenStudy (anonymous):

no worries..could you help me with one more please? x^2-4x-9=0 @03453660

OpenStudy (anonymous):

a)9 b)-4 c)4 d)13

OpenStudy (anonymous):

this is 4

OpenStudy (anonymous):

did you know how to solve this. its a little bit tricky

OpenStudy (anonymous):

ok i can try and figure it out! thank you :)

OpenStudy (anonymous):

you are wlcome